OpenStudy (anonymous):

y = –x + 2 y = 3x – 1 what is the solution of the system

jimthompson5910 (jim_thompson5910):

y = –x + 2 3x - 1 = –x + 2 ... plug in y = 3x - 1 now solve 3x - 1 = –x + 2 for x

jimthompson5910 (jim_thompson5910):

what do you get when you solve for x

OpenStudy (anonymous):

3 right?

jimthompson5910 (jim_thompson5910):

3x-1=-x+2 3x=-x+2+1 3x+x=2+1 4x=2+1 4x=3 x = ???

OpenStudy (anonymous):

0.75

jimthompson5910 (jim_thompson5910):

or x = 3/4

jimthompson5910 (jim_thompson5910):

now use either equation to find y

OpenStudy (anonymous):

can you set it up for me or show me how to set it up?

jimthompson5910 (jim_thompson5910):

y = -x + 2 ... first given equation y = -3/4 + 2 ... plug in x = 3/4 y = ???

OpenStudy (anonymous):

1.25?

jimthompson5910 (jim_thompson5910):

or 5/4

jimthompson5910 (jim_thompson5910):

so x = 3/4, y = 5/4

jimthompson5910 (jim_thompson5910):

in decimal form x = 0.75, y = 1.25

OpenStudy (anonymous):

wow thank you !

jimthompson5910 (jim_thompson5910):

you're welcome